OpenStudy (anonymous):

How do I find the slant height of a cone with a lateral area of 555 pi cm2. the radius is 15. A. 37.0cm2 B. 27.1 C. 22.2 D. 7.7 WILL FRIEND AND GIVE MEDAL

OpenStudy (anonymous):

LA = pi r s

OpenStudy (anonymous):

its the ratio of circumfrences to the original circles area LA = r/s pi s^2 LA = rs pi

OpenStudy (anonymous):

okay can i get a walk though on this.

jimthompson5910 (jim_thompson5910):

L = lateral surface area r = radius s = slant height pi = 3.14 approximately L = pi*r*s

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):